In space far enough from earth’s gravity Superman who had a mass 65kg pushes a space station with 5000 Newton’s of force further away from earth, what will happen to Superman and in what direction.


1
Expert's answer
2021-03-10T17:14:26-0500

Superman get acceleration to the Earth:


"a=\frac{F}{m}\\a=\frac{5000}{65}=76.9\frac{m}{s^2}"